OpenStudy (anonymous):

here is another one I got wrong and do not understand what is the ratio of 28:48 I think it is 20 but I am not sure how to do it

OpenStudy (anonymous):

a ratio is a division, you divide 28 by 48, and that is called simplyfying you are looking for common factors example 44/22=2.2.11/2.11=2

OpenStudy (anonymous):

oh okay sorry I am not good at math these are questions from a test I got these wrong and I want to learn how to do them so I know next time

OpenStudy (anonymous):

well can you factor 28 into primes? do you know what a prime number is?

OpenStudy (anonymous):

a prime number is a number that can only be divided by 1 and itself, ex 17

OpenStudy (anonymous):

example factoring 58, first i see it's even ,so i can divide by 2, i get 58=2*29, but 29 has no divisors(other than 29 and 1) so i am done, 58=2*29

OpenStudy (anonymous):

I believe that the question is to find the easiest representation of the ratio 28:48? In this case, find the largest common denominator, which is done as mivanov says.

OpenStudy (anonymous):

but he must know what a divisor is, to be able to find the divisors of 28 and 48, and after that find the largest common denominator. If he does not know what division is, he can't do it.

OpenStudy (anonymous):

okay well thank you guys mivanov I am a girl and I do know what division is

OpenStudy (anonymous):

sorry if i offended, i am just trying to help. and you have no picture, so i couldn't have known that you're a girl

OpenStudy (anonymous):

no problem

OpenStudy (anonymous):

Well, have you found the answer alli? :) One hint, 14:24 is the same ratio but it is not the most simplified one yet.

OpenStudy (anonymous):

working on it

OpenStudy (anonymous):

okay so 28/48 1